OpenStudy (anonymous):

I need some help, BADLY! Find the solution(s) to the system. Round answers to the nearest hundredth, if necessary. y = - 0.5x^2 + x + 6 y = 3.25

OpenStudy (anonymous):

I'm unsure what the "Y=3.25" is for..

OpenStudy (solomonzelman):

We got one horizontal line, \(\normalsize\color{blue}{ y = 3.25}\) And we got an opening down porabola, \(\normalsize\color{blue}{y = - 0.5x^2 + x \color{red}{-0.5+0.5}+ 6}\) \(\normalsize\color{blue}{y =\underline{ - 0.5x^2 + x \color{red}{-0.5}}\color{red}{+0.5}+ 6}\) \(\normalsize\color{blue}{y = - 0.5(x^2 - 2x + 1)+0.5+6}\) \(\normalsize\color{blue}{y = - 0.5(x - 1)^2+0.5+6}\) \(\normalsize\color{blue}{y = - 0.5(x - 1)^2+6.5}\)

OpenStudy (solomonzelman):

All you have to do, is to find 2 points that are on y=3.25, if there are any.

OpenStudy (anonymous):

Thank you.

OpenStudy (solomonzelman):

Have you found the points ?

OpenStudy (anonymous):

Not yet. I'm trying to figure it out..

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{y = - 0.5(x - 1)^2+6.5}\) \(\normalsize\color{blue}{3.25= - 0.5(x - 1)^2+6.5}\), multiply everything times 4 \(\normalsize\color{blue}{13= -2(x - 1)^2+26}\) \(\normalsize\color{blue}{-13= -2(x - 1)^2}\) \(\normalsize\color{blue}{13= 2(x - 1)^2}\) \(\normalsize\color{blue}{13/2= (x - 1)^2}\) \(\normalsize\color{blue}{± \sqrt{13/2}= (x - 1)}\) \(\normalsize\color{blue}{1± \sqrt{13/2}= x }\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{1± \sqrt{13/2}= x }\) \(\normalsize\color{blue}{1± \sqrt{26/4}= x }\) \(\normalsize\color{blue}{1± \frac{1}{2}\sqrt{26}= x }\)

OpenStudy (solomonzelman):

( ± {1 + ½ √26 } , 3.25 )

OpenStudy (anonymous):

-2, 4

OpenStudy (anonymous):

Right? @SolomonZelman

OpenStudy (solomonzelman):

No, ( ± {1 + ½ √26 } , 3.25 ) ( ± {1 + ½ (5.10) } , 3.25 ) ( ± {1 + 2.05 } , 3.25 ) ( ± 3.05 , 3.25 ) ( 3.05 , 3.25 ) and ( - 3.05 , 3.25 )

OpenStudy (anonymous):

Thank you.